Can someone please answer the following equation:
(40 x a)+(20 x (1.5 x a)) = 600
thanks

Answer by checkley71(8403)

(40a)+(20(1.5a))=600 or 40a+30a=600 or 70a=600 or a=600/70 or a=8.57
proof: (40x8.57)+(20x1.5x8.57)=600 or 343+257=600 600=600.